# Chiral symmetry restoration in QC2D from effective model using the functional renormalization group

### v1

@@ I Introduction
In two-color quantum chromodynamics (QC2D), it is well known that the so-called sign problem in Monte Carlo simulations disappears even in the presence of a nonzero quark-number chemical potential, in contrast to three-color QCD [1]. Hence, first-principles lattice computations are applicable at finite density, therefore, QC2D is regarded as a useful testing ground for elucidating a wide regime of the phase diagram, particularly the cold and dense regime, and has attracted significant attention in recent times [2, 3, 4, 5, 6, 7, 8, 9, 10, 11]. While QC2D does not display the sign problem, it does share a key property with QCD: at low temperatures and/or densities, chiral symmetry is spontaneously broken. To achieve a satisfactory understanding of the strongly interacting matter in QC2D, it is important to identify the similarities and differences in chiral symmetry breaking compared to ordinary QCD. In this paper, our aim is to provide new insights into the order of the phase transition leading to the spontaneous breaking of chiral symmetry for two colors ($N_{c}=2$) and two flavors ($N_{f}=2$) in the zero quark mass limit.
@@
This has already been worked out for $N_{c}=3$, but the order of the transition in the chiral limit remains a subject of debate. According to the traditional scenario based on the $\epsilon$-expansion of renormalization group flows, the transition is first-order for any number of flavors, with a possible exception in the $N_{f}=2$ case, where a strong axial anomaly might induce a second-order transition [13]. However, recent evidence increasingly suggests that the transition may be second-order regardless of the number of flavors, presumably up to the conformal window [14, 15, 16, 17, 18, 19]. Although the corresponding renormalization group fixed point cannot be accessed using the $\epsilon$-expansion, the functional renormalization group (FRG) framework provides clear indications of its existence [20, 21].
No such analyses, including the $\epsilon$-expansion, have been conducted for QC2D so far. Similar to the original $N_{c}=3$ argument, the absence of an infrared-stable fixed point at the critical temperature would suggest that the system cannot exhibit scaling behavior. Therefore, if the transition does indeed exist, it cannot be second order and is presumably first order. Following a similar line of reasoning, in this paper, we aim to investigate the renormalization group flows of the Ginzburg–Landau free energy functional for QC2D with $N_{f}=2$ to search for infrared-stable fixed points that could correspond to a second-order transition. Our approach is twofold: on the one hand, we apply the traditional $\epsilon$-expansion around $d=4$, ultimately extrapolating the $\beta$-functions to $d=3$. On the other hand, we utilize the functional renormalization group (FRG) technique to make predictions directly in $d=3$ without relying on perturbation theory.
The paper is organized as follows. In Sec. II, we define the model and present our ansatz for the free energy. In Sec. III, we outline the procedure for obtaining the scale dependencies (i.e., the $\beta$-functions) of each coupling using the functional renormalization group (FRG) formalism. We first examine the fixed-point structure through the $\epsilon$-expansion and then evaluate the flows directly in $d=3$ using the FRG. Finally, Sec. IV summarizes our findings.
@@ II Effective model
The order parameter in the Ginzburg–Landau (GL) free energy for $N_{c}=2$, $N_{f}=2$ QCD is a $4\times 4$ matrix, belonging to a subset of the $U(4)$ algebra in the following sense [11, 12]:
@@
where $\hat{E}$ is the 4×4$\times$4 ×4 symplectic matrix,
@@
Here $\tau^{0}={\bf 1}$, $\tau^{a=1,2,3}$ are the Pauli matrices, and $D^{4}=\tau^{2}$, $D^{5}=i\tau^{2}$. We stress that, in (1) all parity partners are included; physically $S_{a}$ and $P_{a}$ represent scalar and pseudoscalar mesons while $B_{i}^{\prime}$ and $B$ represent negative-parity and positive-parity diquark baryons, resulting in a total of twelve degrees of freedom. Note that, without all parity partners introduced appropriately, we would found that the RG equations did not close with respect to couplings of the UV free energy.
@@
where the first two lines in the right hand side are fully symmetric under chiral rotations, while the remaining terms break $U_{A}(1)$ transformations  [^1] Throughout this paper, we occasionally use chiral symmetry to refer to Pauli-Gürsey symmetry.. Note that, in contrast to the $N_{c}=3$ system, the standard Kobayashi-Maskawa-’t Hooft determinant is not an independent invariant:
@@
therefore, is excluded from (5). This also means that the only way to incorporate the $U_{A}(1)$ anomaly into the system is through the operator $\sim\,\textrm{Tr}\,[\tilde{\Sigma}\Sigma+\tilde{\Sigma}^{\dagger}\Sigma^{\dagger}]$. That is to say, the form of (5) is the most general expression allowed by symmetry and (perturbative) renormalizability in $d=3$  [^2] In $d=3$, dimensionless couplings correspond to ${\cal O}(\Sigma^{6})$ terms.. We also note that this is the first FRG study that takes into account the $U_{A}(1)$ anomaly appropriately for $N_{c}=2$.
@@
III Renormalization group equations and fixed point analyses
The scale dependent free energy satisfies the Wetterich equation [22]:
@@ III Renormalization group equations and fixed point analyses
Note that $V_{k}$ is considered to be a function of chiral invariants, therefore, when evaluating (12) one differentiates with respect to the invariants first, and then takes the derivatives of the invariants with respect to the fields. After plugging the ansatz (9) into (11), we get, by definition, on the left-hand side the sum of the logarithmic scale derivatives of each coupling, multiplied by their respective operators. The right-hand side must be compatible with this structure, i.e., after taking the inverse and expanding the expression with respect to the components of the background field, the same operator structure must emerge. This allows us to identify, at each order, the corresponding flows of the couplings. Note that one does not need to calculate the inverse matrix of $k^{2}+V_{k}^{\prime\prime}$ in the most general ($2\times 6=12$ component) background field; it is sufficient to consider, at each order, such a background, which allows for the unique separation of each invariant that appears at that given order. We tried several different background fields and always obtained the same results. It is important to emphasize that the structure dictated by (9) is, therefore, indeed inherited by the right-hand side of the flow equation (11), which means that all independent invariants have been properly included in the effective potential.
@@ III.1 $\epsilon$ expansion
The fixed points corresponding to the zeros of Eqs. (16) are summarized in Table I for $\epsilon=1$ ($d\to 3$). Since Eqs. (16) are invariant under the simultaneous sign flips $\bar{a}_{2}\to-\bar{a}_{2}$ and $\bar{a}_{22}\to-\bar{a}_{22}$, we can choose $\bar{a}_{2}<0$. Apart from the trivial Gaussian fixed point, we find an $O(12)$ fixed point where all the anomalous couplings $\bar{a}_{2}$, $\bar{a}_{4}$, $\bar{a}_{11}$, and the quartic coupling $\bar{g}_{2}$ vanish. (Note that $\bar{g}_{2}$ is the coupling that breaks the $O(12)$ invariance of an anomaly-free theory.) Additionally, we find three fixed points (A1, A2, A3) with non-vanishing anomaly, where $\bar{g}_{2}=0$, and two more (A4, A5) with $\bar{g}_{2}\neq 0$, and the anomaly is also present.
@@ III.2 FRG treatment
In this section, we discuss the renormalization group flows and fixed points directly in $d=3$. First, we note that although Eq. (11) may have been used to derive the results of the $\epsilon$ expansion, in principle, the FRG treatment is not restricted to working close to $d=4$. Eq. (11) is non-perturbative in the sense that there is no need to assume the existence of any small parameter. It follows from Eq. (10) with the ansatz given by Eq. (8), which serves as the leading order approximation of the derivative expansion. The latter has been shown to converge when using the optimal regulator defined above Eq. (11) [28], so evaluating the flows directly at $d=3$ without expanding in any small parameter is justified.
The calculation of each $\beta$ function follows the procedure outlined at the beginning of this section. The main differences between the direct $d=3$ calculation and its $d=4-\epsilon$ counterpart are as follows. In the $d=4-\epsilon$ case, the flows are relatively simple because we are working at leading order in $\epsilon$, whereas in $d=3$ one cannot make such simplifications, and the expressions for the $\beta$ functions are much more involved. Additionally, since in $d=3$ the marginal interactions are of ${\cal O}(\Sigma^{6})$, the number of couplings in the model is twelve [as shown in Eq. (8)]. We do not attempt to present the expression for each flow here, but the corresponding formulas can be found in the supplementary material.
@@
The resulting fixed point structure is shown in Table II. This structure is quite similar to that of Table I, with two main differences. First, the A2 and A5 fixed points split into two, resulting in the creation of A2b and A5b. Second, and more importantly, we observe the emergence of two completely new, anomaly-free fixed points (AF1 and AF2), which were not accessible in the $\epsilon$ expansion. This resembles the $N_{c}=3$ case, where such a fixed point is infrared stable at the critical point if the $U_{A}(1)$ anomaly is absent [20, 21]. This is also the case in our current $N_{c}=2$ scenario. If the $U_{A}(1)$ anomaly is present at the critical point, for both AF1 and AF2 the number of relevant directions is three, leaving no possibility for an attractive fixed point at the transition point. However, if the $U_{A}(1)$ anomaly disappears at the critical point, the anomalous directions are no longer present, reducing the number of relevant directions for the AF fixed points to one. In this case, both of the AF fixed points can correspond to a continuous phase transition at finite temperature. Note, however, that, the the sign of the $g_{2}$ coupling (and accordingly, that of $c_{1}$) is different for them. One notes that AF1 realizes a potential that is bounded from below, however, AF2 does not (see next subsection). That is, for our purposes, it is expected that the AF1 fixed point can be a candidate for a second order chiral phase transition, even though by improving the approximation, it might turn out that AF2 also corresponds to a stable potential, therefore, one should not entirely rule out its significance. The flow chart corresponding to all non-anomalous fixed points is sketched in Fig. 1.
@@
We also note that the analysis presented in the previous subsection regarding the $O(6)$ fixed point still holds within the FRG approach. The corresponding fixed point is shown in the last line of Table II. As a result, we identify two potential candidates for fixed points that could be responsible for a continuous phase transition, which aligns with the situation in the $N_{c}=3$, $N_{f}=2$ system. As discussed in the introduction, for $N_{c}=3$, there is increasing evidence suggesting that the chiral transition is second-order for zero quark masses. It would be interesting to determine whether the $N_{c}=2$ system exhibits the same property. Our current renormalization group analyses indicate that the transition could indeed be second-order.
III.3 Stability of the fixed point potentials
@@ IV Summary
In this paper, we investigated whether the effective Ginzburg–Landau (GL) free energy functional for $N_{c}=2$ quantum chromodynamics (QC2D) with $N_{f}=2$ flavors can exhibit fixed points corresponding to a finite-temperature second-order phase transition. The free energy for $N_{c}=2$ differs from the $N_{c}=3$ case in that the theory displays an SU(4) Pauli-Gürsey symmetry instead of SU(2)$\times$×SU(2). After constructing the most general renormalizable free energy, we calculated the $\beta$ functions of the couplings both in the $\epsilon$ expansion around $d=4$ and using the functional renormalization group (FRG) approach, which allows for direct evaluation of scale dependencies in $d=3$.
Our findings using the $\epsilon$ expansion indicate that, despite the more intricate fixed point structure found through for $N_{c}=2$ compared to $N_{c}=3$, there is no infrared stable fixed point at the critical temperature for finite axial anomaly couplings, meaning no second-order transition can be associated. However, for an infinitely strong anomaly, an $O(6)$ fixed point, analogous to the $O(4)$ fixed point in the $N_{c}=3$ case, emerges and is infrared stable at the critical point.
The FRG method predicts a much richer structure. In particular, we identified the existence of two fixed points (labeled AF1 and AF2), where all anomalous couplings vanish. These fixed points are not infrared stable in a theory that breaks $U_{A}(1)$; however, in a GL functional that preserves $U_{A}(1)$, the stability analysis changes, and both fixed points are found to be infrared stable at the critical point. While AF2 corresponds to an unphysical potential in our approximation, the AF1 fixed point can be considered a candidate to describe a finite temperature second order phase transition.
@@
Exploring the chiral phase transition at finite temperature in QC2D may provide valuable insights into $N_{c}$ dependencies. Although current QC2D lattice simulations employ relatively heavy quark masses, it would be interesting to determine whether future simulations can detect any signs of a continuous phase transition in the chiral limit. Also, our analysis could be extended by investigating the role of anomalous dimensions at each fixed point and examining the influence of non-renormalizable interactions in the GL functional, assessing their impact on the fixed-point structure. Additionally, the FRG method enables the determination of fully non-perturbative fixed-point potentials without assuming any specific functional form, potentially providing further evidence for the existence of infrared-stable fixed points. These aspects will be explored in future work and reported elsewhere.
Acknowledgements
This work was supported by the Hungarian National Research, Development, and Innovation Fund under Project No. FK142594. G.F. would like to express his gratitude for the warm hospitality of the Kobayashi-Maskawa Institute for the Origin of Particles and the Universe at Nagoya University, where this research project was initiated. D.S. was supported by the JSPS KAKENHI Grant No. 23K03377 and No. 23H05439. D.S. expresses gratitude to the Young Researchers Overseas Program in Science of Nagoya University, which supported his stay at Eötvös Loránd University.

### v2

@@ I Introduction
In two-color quantum chromodynamics (QC2D), it is well known that the so-called sign problem in Monte Carlo simulations disappears even in the presence of a nonzero quark-number chemical potential, in contrast to three-color QCD [1]. Hence, first-principles lattice computations are applicable at finite density: therefore, QC2D is regarded as a useful testing ground for elucidating a wide regime of the phase diagram, particularly the cold and dense regime, and has attracted significant attention in recent times [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12]. While QC2D does not display the sign problem, it does share a key property with QCD: at low temperatures and/or densities, chiral symmetry is spontaneously broken. To achieve a satisfactory understanding of the strongly interacting matter in QC2D, it is important to identify the similarities and differences in chiral symmetry breaking compared to ordinary QCD. In this paper, our aim is to provide new insights into the order of the phase transition leading to the spontaneous breaking of chiral symmetry for two colors ($N_{c}=2$) and two flavors ($N_{f}=2$) in the zero quark mass limit.
@@
This has already been worked out for $N_{c}=3$, but the order of the transition in the chiral limit remains a subject of debate. According to the traditional scenario based on the $\epsilon$ expansion of renormalization group flows, the transition is first order for any number of flavors, with a possible exception in the $N_{f}=2$ case, where a strong axial anomaly might induce a second-order transition [14]. However, recent evidence increasingly suggests that the transition may be second order regardless of the number of flavors, presumably up to the conformal window [15, 16, 17, 18, 19, 20]. Although the corresponding renormalization group fixed point cannot be accessed using the $\epsilon$ expansion, the functional renormalization group (FRG) framework provides clear indications of its existence [21, 22].
No such analyses, including the $\epsilon$ expansion, have been conducted for QC2D so far. Similar to the original $N_{c}=3$ argument, the absence of an infrared-stable fixed point at the critical temperature would suggest that the system cannot exhibit scaling behavior. Therefore, if the transition does indeed exist, it cannot be second order and is presumably first order. Following a similar line of reasoning, in this paper, we aim to investigate the renormalization group flows of the GL free energy functional for QC2D with $N_{f}=2$ to search for infrared-stable fixed points that could correspond to a second-order transition. Our approach is twofold: on the one hand, we apply the traditional $\epsilon$ expansion around $d=4$, ultimately extrapolating the $\beta$ functions to $d=3$. On the other hand, we utilize the FRG technique to make predictions directly in $d=3$ without relying on perturbation theory.
The paper is organized as follows. In Sec. II, we define the model and present our ansatz for the free energy. In Sec. III, we outline the procedure for obtaining the scale dependencies (i.e., the $\beta$ functions) of each coupling using the FRG formalism. We first examine the fixed-point structure through the $\epsilon$ expansion and then evaluate the flows directly in $d=3$ using the FRG. Finally, Sec. IV summarizes our findings.
@@ II Effective model
The order parameter in the GL free energy for $N_{c}=2$, $N_{f}=2$ QCD is a $4\times 4$ matrix, belonging to a subset of the $U(4)$ algebra in the following sense [12, 13]:
@@
where $\hat{E}$ is the $4\times 4$ symplectic matrix,
@@
Here $\tau^{0}={\bf 1}$, $\tau^{a=1,2,3}$ are the Pauli matrices, and $D^{4}=\tau^{2}$, $D^{5}=i\tau^{2}$. We stress that, in (1) all parity partners are included; physically $S_{a}$ and $P_{a}$ represent scalar and pseudoscalar mesons while $B_{i}^{\prime}$ and $B$ represent negative-parity and positive-parity diquark baryons, resulting in a total of 12 degrees of freedom. Note that without all parity partners introduced appropriately, we found that the RG equations did not close with respect to couplings of the UV free energy.
@@
where the first two lines in the right-hand side are fully symmetric under chiral rotations, while the remaining terms break $U_{A}(1)$ transformations [^1] Throughout this paper, we occasionally use chiral symmetry to refer to Pauli-Gürsey symmetry.. Note that in contrast to the $N_{c}=3$ system, the standard Kobayashi-Maskawa-’t Hooft determinant is not an independent invariant:
@@
[…]
@@
III Renormalization group equations and fixed-point analyses
The scale-dependent free energy satisfies the Wetterich equation [23]:
@@ III Renormalization group equations and fixed-point analyses
Note that $V_{k}$ is considered to be a function of chiral invariants; therefore, when evaluating (12) one differentiates with respect to the invariants first, and then takes the derivatives of the invariants with respect to the fields. After plugging the ansatz (9) into (11), we get, by definition, on the left-hand side the sum of the logarithmic scale derivatives of each coupling, multiplied by their respective operators. The right-hand side must be compatible with this structure, i.e., after taking the inverse and expanding the expression with respect to the components of the background field, the same operator structure must emerge. This allows us to identify, at each order, the corresponding flows of the couplings. Note that one does not need to calculate the inverse matrix of $k^{2}+V_{k}^{\prime\prime}$ in the most general ($2\times 6=12$ component) background field; it is sufficient to consider, at each order, such a background, which allows for the unique separation of each invariant that appears at that given order. We tried several different background fields and always obtained the same results. It is important to emphasize that the structure dictated by (9) is, therefore, indeed inherited by the right-hand side of the flow equation (11), which means that all independent invariants have been properly included in the effective potential.
@@ III.1 $\epsilon$ expansion
The fixed points corresponding to the zeros of Eqs. (16) are summarized in Table I for $\epsilon=1$ ($d\to 3$). Since Eqs. (16) are invariant under the simultaneous sign flips $\bar{a}_{2}\to-\bar{a}_{2}$ and $\bar{a}_{22}\to-\bar{a}_{22}$, we can choose $\bar{a}_{2}<0$. Apart from the trivial Gaussian fixed point, we find an $O(12)$ fixed point where all the anomalous couplings $\bar{a}_{2}$, $\bar{a}_{4}$, $\bar{a}_{22}$, and the quartic coupling $\bar{g}_{2}$ vanish. [Note that $\bar{g}_{2}$ is the coupling that breaks the $O(12)$ invariance of an anomaly-free theory.] Additionally, we find three fixed points (A1, A2, A3) with nonvanishing anomaly, where $\bar{g}_{2}=0$, and two more (A4, A5) with $\bar{g}_{2}\neq 0$, and the anomaly is also present.
@@ III.2 FRG treatment
In this section, we discuss the renormalization group flows and fixed points directly in $d=3$. First, we note that although Eq. (11) may have been used to derive the results of the $\epsilon$ expansion, in principle, the FRG treatment is not restricted to working close to $d=4$. Equation (11) is nonperturbative in the sense that there is no need to assume the existence of any small parameter. It follows from Eq. (10) with the ansatz given by Eq. (8), which serves as the leading-order approximation of the derivative expansion. The latter has been shown to converge when using the optimal regulator defined above Eq. (11) [29], so evaluating the flows directly at $d=3$ without expanding in any small parameter is justified.
The calculation of each $\beta$ function follows the procedure outlined at the beginning of this section. The main differences between the direct $d=3$ calculation and its $d=4-\epsilon$ counterpart are as follows. In the $d=4-\epsilon$ case, the flows are relatively simple because we are working at leading order in $\epsilon$, whereas in $d=3$ one cannot make such simplifications, and the expressions for the $\beta$ functions are much more involved. Additionally, since in $d=3$ the marginal interactions are of ${\cal O}(\Sigma^{6})$, the number of couplings in the model is 12 [as shown in Eq. (8)]. We do not attempt to present the expression for each flow here, but the corresponding formulas can be found in the Supplemental Material.
@@
The resulting fixed-point structure is shown in Table II. This structure is quite similar to that of Table I, with two main differences. First, the A2 and A5 fixed points split into two, resulting in the creation of A2b and A5b. Second, and more importantly, we observe the emergence of two completely new, anomaly-free fixed points (AF1 and AF2), which were not accessible in the $\epsilon$ expansion. This resembles the $N_{c}=3$ case, where such a fixed point is infrared stable at the critical point if the $U_{A}(1)$ anomaly is absent [21, 22]. This is also the case in our current $N_{c}=2$ scenario. If the $U_{A}(1)$ anomaly is present at the critical point, for both AF1 and AF2 the number of relevant directions is three, leaving no possibility for an attractive fixed point at the transition point. However, if the $U_{A}(1)$ anomaly disappears at the critical point, the anomalous directions are no longer present, reducing the number of relevant directions for the AF fixed points to one. In this case, both of the AF fixed points can correspond to a continuous phase transition at finite temperature. Note, however, that, the sign of the $g_{2}$ coupling (and accordingly, that of $c_{1}$) is different for them. One notes that AF1 realizes a potential that is bounded from below, however, AF2 does not (see next subsection). That is, for our purposes, it is expected that the AF1 fixed point can be a candidate for a second-order chiral phase transition, even though by improving the approximation, it might turn out that AF2 also corresponds to a stable potential; therefore, one should not entirely rule out its significance. The flow chart corresponding to all nonanomalous fixed points is sketched in Fig. 1.
@@
We also note that the analysis presented in the previous subsection regarding the $O(6)$ fixed point still holds within the FRG approach. The corresponding fixed point is shown in the last line of Table II. As a result, we identify two potential candidates for fixed points that could be responsible for a continuous phase transition, which aligns with the situation in the $N_{c}=3$, $N_{f}=2$ system. As discussed in the Introduction, for $N_{c}=3$, there is increasing evidence suggesting that the chiral transition is second order for zero quark masses. It would be interesting to determine whether the $N_{c}=2$ system exhibits the same property. Our current renormalization group analyses indicate that the transition could indeed be second order.
III.3 Stability of the fixed-point potentials
@@ IV Summary
In this paper, we investigated whether the effective GL free energy functional for QC2D with $N_{f}=2$ flavors can exhibit fixed points corresponding to a finite-temperature second-order phase transition. The free energy for $N_{c}=2$ differs from the $N_{c}=3$ case in that the theory displays an SU(4) Pauli-Gürsey symmetry instead of SU(2)$\times$×SU(2). After constructing the most general renormalizable free energy, we calculated the $\beta$ functions of the couplings both in the $\epsilon$ expansion around $d=4$ and using the FRG approach, which allows for direct evaluation of scale dependencies in $d=3$.
Our findings using the $\epsilon$ expansion indicate that, despite the more intricate fixed-point structure found for $N_{c}=2$ compared to $N_{c}=3$, there is no infrared stable fixed point at the critical temperature for finite axial anomaly couplings, meaning no second-order transition can be associated. However, for an infinitely strong anomaly, an $O(6)$ fixed point, analogous to the $O(4)$ fixed point in the $N_{c}=3$ case, emerges and is infrared stable at the critical point.
The FRG method predicts a much richer structure. In particular, we identified the existence of two fixed points (labeled AF1 and AF2), where all anomalous couplings vanish. These fixed points are not infrared stable in a theory that breaks $U_{A}(1)$; however, in a GL functional that preserves $U_{A}(1)$, the stability analysis changes, and both fixed points are found to be infrared stable at the critical point. While AF2 corresponds to an unphysical potential in our approximation, the AF1 fixed point can be considered a candidate to describe a finite temperature second-order phase transition.
@@
Exploring the chiral phase transition at finite temperature in QC2D may provide valuable insights into $N_{c}$ dependencies. Although current QC2D lattice simulations employ relatively heavy quark masses, it would be interesting to determine whether future simulations can detect any signs of a continuous phase transition in the chiral limit. Also, our analysis could be extended by investigating the role of anomalous dimensions at each fixed point and examining the influence of nonrenormalizable interactions in the GL functional, assessing their impact on the fixed-point structure. Additionally, the FRG method enables the determination of fully nonperturbative fixed-point potentials without assuming any specific functional form, potentially providing further evidence for the existence of infrared-stable fixed points. These aspects will be explored in future work and reported elsewhere.
Acknowledgments
This work was supported by the Hungarian National Research, Development, and Innovation Fund under Project No. FK142594. G.F. would like to express his gratitude for the warm hospitality of the Kobayashi-Maskawa Institute for the Origin of Particles and the Universe at Nagoya University, where this research project was initiated. D.S. was supported by the JSPS KAKENHI Grants No. 23K03377 and No. 23H05439. D.S. expresses gratitude to the Young Researchers Overseas Program in Science of Nagoya University, which supported his stay at Eötvös Loránd University.